OpenStudy (anonymous):

Solve the equation. 7(2x) = 23 - 6/2x

OpenStudy (anonymous):

\[14X ^{2} - 23X +3 = 0\]\[X _{1}=3/2 ; X _{2}=1/7\]

OpenStudy (anonymous):

how did you get 14x^2 - 23x + 3 = 0 ?

OpenStudy (anonymous):